A student has a job that leaves her with $300 per month in disposable

income. She decides that she will use the money to buy a car. Before

looking for a car, she arranges a 100% loan whose terms are $300 per

month for 48 months at 9% nominal annual interest. What is the

maximum car purchase price that she can afford with her loan?


1
Expert's answer
2021-10-01T16:55:01-0400

Present value factor or present value interest factor is widely utilized to forecast the present or current worth or value of the money at a particular point in the future. It helps to compare the worthiness of cash in different time periods.

Maximum price=Payment per period"×\frac{1-(1+r)^{-n}}{r}"

Substitute r with "\frac{9\%}{12}=0.0075" and n with 48

Maximum price=Payment per period"×\frac{1-(1+0.075)^{-48}}{0.0075}=\$12055"